OpenStudy (anonymous):

what is the solution set of x=2(square root)2x-3

OpenStudy (jdoe0001):

\(\bf x = \sqrt{2x-3}\implies x^2=2x-3\implies x^2-2x+3=0\) factor it like you'd any quadratic